OpenStudy (anonymous):

Find the equation of a circle that has its center at (6, -4) and has a radius of 3. (x+ 6)2+ (y- 4)2= 9 (x- 6)2+ (y- 4)2= 9 (x+ 6)2+ (y+ 4)2= 9 (x- 6)2+ (y+ 4)2= 9

OpenStudy (lgbasallote):

hmm so what do you think is the answer here?

OpenStudy (anonymous):

I think it is the first one

OpenStudy (lgbasallote):

why do you think so?

OpenStudy (anonymous):

because the center is (6,-4) and in the first equation (x+ 6)2+ (y- 4)2= 9

OpenStudy (lgbasallote):

but the equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\] right?

OpenStudy (lgbasallote):

h is the x-coordinate of the center k is the y-coordinate of the center

OpenStudy (anonymous):

hold on i am going to use the equation quick

OpenStudy (lgbasallote):

sure :)

OpenStudy (anonymous):

I am not sure if I did it correct still but, i actually got the last one

OpenStudy (lgbasallote):

ahh that's right! congrats

OpenStudy (anonymous):

thank you:D

OpenStudy (lgbasallote):

<tips aht>

OpenStudy (lgbasallote):

hat*